OpenStudy (anonymous):

Rectangle ADEF and isosceles trapezoid FBCE both have a base of 12. The lower base angles of the trapezoid are 45° and the legs of the trapezoid are 6. How long is BC? Round your answer to the nearest tenth.

OpenStudy (anonymous):

2. Draw FBCE with BC at the top, at FE as the base. Draw lines straight down from B and C to FE, to form triangles with angles, 45, 45, 90. This is a 1:1:sqrt(2) triangle the length of each of the sections is 6 / sqrt(2) = 3 * sqrt (2) There are 2 of the triangles. 12 - 2 * (3 * sqrt (2)) = 12 - 6 sqrt(2) = 12 - 8.485 = 3.5